OpenStudy (anonymous):

A bag of marbles contains only purple and yellow marbles. The bag has 3 times as many yellow marbles as purple. The bag contains 164 total marbles. Write an equation or system of equations to find out how many of each color of marble is in the bag.

OpenStudy (anonymous):

x+3x=164 x=41 there are 41purple marbles, and 123 yellow marbles

OpenStudy (anonymous):

Thank You (:

OpenStudy (anonymous):

Can you explain to me how you got that?

OpenStudy (anonymous):

you use x+3x=164 because it is telling you that you have an original amount of marbles (x) and the other color is 3 times that original amount (3x). x+3x=164 combine like terms to get, 4x=164 divide both sides by 4 to get x alone, x=41 so now that you know there are 41 purple marbles, you can solve for the yellow marbles because you know it is equal to 3x. 3x (or 3 times 41) =123

OpenStudy (anonymous):

and finally, you can check your answer by adding 123+41, to get 164

OpenStudy (anonymous):

Thanks!